OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

2x+6 can be equal to either +4 or -4 when 2x+6=4 x=-1 when 2x+6=-4 x=-5 hence x=-1,-5

OpenStudy (anonymous):

thank you for your help :)